OpenStudy (anonymous):

Bruce had an EKG to measure his heartbeat rate. After conversion, the function produced could be modeled by a cosine function, and the wave produced a maximum of 4, minimum of −2, and period of pi over 2. Which of the following functions could represent Bruce's EKG read-out?

OpenStudy (anonymous):

f(x) = 4 cos pi over 2x − 2 f(x) = 3 cos 4x + 1 f(x) = 3 cos pi over 2x + 1 f(x) = 4 cos 4x − 2

OpenStudy (anonymous):

f(x) = A cos (Bx + C) + D A = amplitude = ½|max - min| B = (2π)/period C = phase shift = 0 for this problem D = vertical shift = max - amplitude

OpenStudy (anonymous):

A and D are wrong right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

The coefficient of x is 2pi/period, so it should be \[2 \pi \div \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

oh so its 4

OpenStudy (anonymous):

so 4x

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i see now its B thank you very much

OpenStudy (anonymous):

you're welcome